OpenStudy (fernandez555):

f(x)=x+6 and g(x) = x^4 what is G(F(x))??

OpenStudy (squidnnneyyy):

ok first, solve for f(x)

rebeccaxhawaii (rebeccaxhawaii):

Plug in f(x) for g(x) - so g(f(x))=(x+6)^4

OpenStudy (fernandez555):

distribution? is that all?

rebeccaxhawaii (rebeccaxhawaii):

\[(x+6)^4=(x+6)(x+6)(x+6)(x+6)\]

OpenStudy (fernandez555):

so, in another scenario. for f(x)=6x and g(x) =4x+1, to find (fog)(x) new equation would be 24x+6

OpenStudy (mathmale):

This problem is quite different from the previous one. In that one, you simply added together functions f(x) and g(x). This problem, on the other hand, is called a "composite function." f(g(x)), read "f of g of x," is evaluated by replacing both x in f(x) = x+6 by g(x). Thus, you get f(g(x))=( ? ) + 6. Please finish this work. f(g(x)) = ?

OpenStudy (fernandez555):

24x

OpenStudy (mathmale):

But you are to find g(f(x)). The method remains the same: Write out g(x). Then replace each appearance of x with f(x). What is f(x) equal to? Please be sure to label your results. '24x' has not been labeled.